Question 4 a. Suppose a firm enters a fixed for floating interest rate swap with a swap dealer. With example, illustrate the cash flow that will occur as a result o the swap (13 marks) Sabah Top Berhad currently has 330,000 shares outstanding with market value of RM64 per share. Assuming no market imperfections or tax effect exist, how much is the share price after: b. i. Sabah Top has a five-for-three share spilt? i. Sabah Top has a 15 percent share dividend? ili. Sabah Top has a RM3 cash share dividend? iv. Sabah Top has a four-for-seven reverse share split? (3 marks) (3 marks) (3 marks) (3 marks TOTAL 25 MARK
Answer #1

Q4

B

The solution is provided below for your reference

i

=64*3/5=38.4

ii

=64/1.15=55.65217

iii

=64-3=61

iv

=64*7/4=112